OpenStudy (anonymous):

I know the answer to my problem I just need someone to explain to me how it's done please wait while I draw my problem Ty :)

OpenStudy (anonymous):

Logbasex 1000= 0.75

OpenStudy (anonymous):

Is this logarithms? :)

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

@AccessDenied @Brittni0605

OpenStudy (anonymous):

answer is 10^4=10000

OpenStudy (anonymous):

Ik but how did u get that answer

OpenStudy (anonymous):

logbasex 1000=3*logbasex 10 3* logbasex 10= 0.75 logbasex 10= 0.25 logbase10 x= 4 x=10^4 do you understand??

OpenStudy (anonymous):

Not rlly. How did u get 10 and 0.25

OpenStudy (anonymous):

logbasex 1000=logbasex (10^3)=3*logbasex 10 eg: logbasex (y^z)=z*logbasex (y) and then 0.75/3=0.25 do you understand??

OpenStudy (anonymous):

Yes I get it know thank you soo much tht is gonna help me on my test :D GBU

OpenStudy (anonymous):

welcome :)